OpenStudy (anonymous):

Solve the quadratic equation x^2 - x = 30 by completing the square. A. {-5, 6} B. {-5, -6} C. {4, 5} D. {-4, -5}

OpenStudy (anonymous):

@ucreategames

OpenStudy (johnweldon1993):

You need to take the coefficient of the 'x' term (-1 here) divide it by 2....then square the result....you then add this to both sides of the equation... so -1/2 = -1/2 (-1/2)^2 = 1/4 Added to both sides x^2 - x + 1/4 = 30 + 1/4 x^2 - x + 1/4 = 121/4 Now turn the left side into a result of squares... (x - (1/2))^2 = 121/4 Take the square root of both sides of the equation x - (1/2) = √121/4 x - (1/2) = +/- 11/2 Break into 2 equations... x - (1/2) = 11/2 x - (1/2) = -11/2 Add 1/2 to both sides of each equation x = 12/2 = 6 and x = -10/2 = -5 (6 and -5) are your answers

OpenStudy (anonymous):

true, but thats not one of the options

OpenStudy (johnweldon1993):

...look at the first choice...

OpenStudy (anonymous):

true i was trippen, lol thank you !!

OpenStudy (johnweldon1993):

Lol no prob!